# Deterministic Bethe state preparation

### v2

@@
Deterministic Bethe state preparation
David Raveh111dxr921@miami.edu and  Rafael I. Nepomechie222nepomechie@miami.edu
Physics Department, PO Box 248046
University of Miami, Coral Gables, FL 33124 USA
@@ 1 Introduction
After [19], an algorithm for preparing Bethe states of the spin-1/2 closed XXZ chain based on coordinate Bethe ansatz was formulated [20]. This algorithm, which was later generalized to the open chain [21], was restricted to real Bethe roots, required ancillary qubits, and was probabilistic. Moreover, the success probability was shown to tend super-exponentially to zero with the number of down-spins [22]. A different approach, based instead on algebraic Bethe ansatz, was subsequently developed for the closed XXZ chain [23]. The latter algorithm was not limited to real Bethe roots, and was deterministic; however, it required performing QR decompositions of matrices whose size scale exponentially with the number of down-spins. Analytic formulae for the unitaries of [23] were proposed in [24].
We present here a new algorithm for preparing arbitrary $L$-qubit $U(1)$-eigenstates, including exact Bethe states of the spin-1/2 XXZ chain with either open or closed boundary conditions. Like the algorithms [20, 21], it is based on coordinate Bethe ansatz; however, it is not limited to real Bethe roots, does not require ancillary qubits, and it is deterministic. Moreover, unlike [23], our algorithm is explicit, and does not make use of QR decompositions. For a state with $M$ down-spins in a chain of length $L$, the circuit uses only $L$ qubits, and has size and depth $\mathcal{O}\left(\binom{L}{M}\right)$. Our algorithm is inspired by an efficient recursive algorithm[25, 26, 27, 28] for preparing Dicke states [29]. Code in Qiskit [30] implementing this algorithm is available as Supplementary Material.
@@ 2 Recursion
where the sum is over the set $P(L,M)$ of all permutations $w$ with $L-M$ zeros (up-spins) and $M$ ones (down-spins), which we think of as $\binom{L}{M}$ strings of length $L$. Moreover, the coefficients $f(w)$ are known complex numbers, which depend on the corresponding $M$ Bethe roots that we assume are known, see Appendix A. Here $F(\{\})$ denotes the normalization
@@
and $|0011\rangle=|0\rangle\otimes|0\rangle\otimes|1\rangle\otimes|1\rangle$, etc., with computational basis states $|0\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $|1\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$. The Bethe states are not the only example of states that take the form (2.1); in fact, the eigenstates of any Hamiltonian with $U(1)$ symmetry take this form. [^1] Indeed, since the states $|B^{L}_{M}\rangle$ form an orthonormal basis, any state $|\phi\rangle$ can be expressed as $|\phi\,\rangle=\sum_{M=0}^{L}C_{M}\,|B^{L}_{M}\rangle$. Hence, if $|\phi\rangle$ is an eigenstate of $S^{z}$ with eigenvalue $\alpha$, then $\alpha\,|\phi\rangle=S^{z}|\phi\rangle=\sum_{M=0}^{L}\,(L/2-M)\,C_{M}\,|B^{L}_{M}\rangle$. This implies $\alpha\,C_{M}=(L/2-M)\,C_{M}$, so either $\alpha=L/2-M$ or $C_{M}=0$. In particular, $C_{M}$ is non-zero for exactly one value of $M$, so $|\phi\rangle=|B^{L}_{M}\rangle$. Although we have in mind values $f(w)$ dictated by Bethe ansatz, and we shall refer to $|B^{L}_{M}\rangle$ as Bethe states, our algorithm is actually more general, as it prepares the state $|B^{L}_{M}\rangle$ for arbitrary complex numbers $f(w)$. In particular, for the case $f(w)=1$, $|B^{L}_{M}\rangle$ reduces to the Dicke state $|D^{L}_{M}\rangle$, whose preparation was shown to have an efficient recursive algorithm [25].
@@
That is, we are restricting the sum in (2.4) to permutations $w$ of the form $w=ab$, where $ab$ denotes the concatenation of the strings $a$ and $b$, so $|w\rangle=|a\rangle\otimes|b\rangle$. The state (2.4) is thus the “$b$-tail” of the Bethe state (2.1). Note that when $b=\{\}$ is the empty string, $F(b)$ reduces to the normalization given by (2.2). For example, for the state $|B^{4}_{2}\rangle$ (2.3) with $b=0$, we have
@@ 3 Algorithm
It follows from (2.7) that $U_{m}$ has the recursive property
@@
It thus remains to construct a circuit $W_{m}$ that performs (3.5) for all $l$ and $b$. This can be achieved using an operator $\textup{I}_{m,l}$ independent of $b$ with the following properties:
@@
for all $l$. In view of the bounds (3.2), we conclude that the operator $W_{m}$, independent of $l$, can be written as the following ordered product of these I operators:
@@
Figure 1: Circuit diagrams for $\textup{I}_{m,l}$, with $\mathcal{U}(m,l)$ defined as a product of $u$-gates (3.12). Additional controls on wires 00 to $L-m-1$ (represented by a rectangle with 3 bullets) are placed on each $u(m,l,b)$ at the locations where $b$ contains a one.
The circuit diagram for an operator $\textup{I}_{m,l}$ with the properties (3.8) and (3.9) is displayed in Fig. 1. There, $\mathcal{U}(m,l)$ denotes the product
@@
Each $u$-gate acts on wire $L-m$, and has controls on the wires $L-m+l$ and $L-m+l-1$ (the latter of which is present if and only if $l>1$).444 We follow Qiskit conventions, where e.g. $|110\rangle$ corresponds to $|0\rangle$ on wire 00 and $|1\rangle$ on wires $1$ and $2$. Additionally, because the angles depend on $b$, additional controls are placed on wires 00 to $L-m-1$ at the locations where the string $b$ takes the value $1$, to differentiate between different $b$ [^5] This method has been used before to generate states with many free parameters, see e.g. [31]. For example, for $L=4,\,M=2$, $u(m=2,l=1,b=01)$ has controls on the wires $3$ and 00, whereas $u(m=2,l=1,b=10)$ has controls on the wires $3$ and $1$. For the case in which $b$ in (3.12) is unique, these additional controls can be omitted. Complete circuit diagrams for examples with $L=4,\,M=2$ and $L=6,\,M=3$ are displayed in Figs. 2 and 3 in Appendix B. An implementation of this algorithm using Qiskit [30] is available as Supplementary Material.
@@
$M+2M(L-M)+{L\choose M}-1=\mathcal{O}\left(\binom{L}{M}\right)\,.$ (3.17)
For the central case $M=L/2$, we note the asymptotic approximation (see e.g. [32], p. 35)
@@
Besides the preparation of the initial state, no two gates in this circuit can run in parallel, so the depth is $\mathcal{O}\left(\binom{L}{M}\right)$ as well. Further, we see that the number of CNOT-gates (3.15) and $u$-gates (3.16) are symmetric under $M\to L-M$, which implies that there is no advantage to using the dual symmetry $|B^{L}_{M}\rangle=X^{\otimes L}\,|B^{L}_{L-M}\rangle$.
@@ 4 Discussion
We have formulated an algorithm for exactly preparing on a quantum computer an $L$-qubit $U(1)$-eigenstate consisting of an arbitrary superposition of computational basis states with $M$ ones (down-spins) and $L-M$ zeros (up-spins) (2.1). The circuit is ideally suited for models solved by Bethe ansatz, which have available general recipes for the values $f(w)$ corresponding to the models’ eigenstates for arbitrary values of $L$ and $M$ in terms of Bethe roots $\{k_{1},...,k_{M}\}$. However, these coefficients $f(w)$, of which there are (${L\choose M}$ many, are each expressed as a sum over $M!$ terms for the closed chain (A.5), and $2^{M}M!$ terms for the open chain (A.12), and are thus expensive to compute classically for large $M$. (For eigenstates of non-integrable models, the $f(w)$ values are generally not known.) The circuit size is of the same order as the number of free parameters $f(w)$, suggesting that the circuit is optimal for given values of $L$ and $M$. Nevertheless, for $M=L/2$, the circuit size and depth scale exponentially with $L$. As noted in the Introduction, these states could be used to compute correlation functions [17, 18].
@@
We have assumed here, as in [20, 21, 22, 23, 24], that the Bethe roots for the desired state are known. Real Bethe roots can readily be determined classically, even for values $L\sim 10^{3}$, by iteration (see, e.g. [34] and references therein). However, complex Bethe roots are generally difficult to find classically. Following up on an idea suggested in [19], it may be possible to implement a hybrid classical-quantum approach (see e.g. [7] and references therein) for determining such complex Bethe roots by using our Bethe states as trial states and the unknown Bethe roots as variational parameters.
@@ A.1 Closed periodic XXZ chain
$f(w)=\sum_{\mathbb{P}}\varepsilon_{\mathbb{P}}\,A(k_{1},\cdots,k_{M})\,e^{i\sum_{j=1}^{M}k_{j}x_{j}}\,,$ (A.5)
where the sum is over all permutations $\mathbb{P}$ of $k_{1},\ldots,k_{M}$, and $\varepsilon_{\mathbb{P}}=\pm 1$ changes sign at each such permutation. Moreover,
@@ A.2 Open XXZ chain
$f(w)=\sum_{\mathbb{P}}\varepsilon_{\mathbb{P}}\,A(k_{1},\ldots,k_{M})\,e^{i\sum_{j=1}^{M}k_{j}x_{j}}\,,$ (A.12)
where now the sum is over all permutations and negations of $k_{1},\ldots,k_{M}$, and $\varepsilon_{\mathbb{P}}=\pm 1$ changes sign at each such mutation. Moreover, now

### v3

@@
Abstract
@@ 1 Introduction
After [19], an algorithm for preparing Bethe states of the spin-1/2 closed XXZ chain based on the coordinate Bethe ansatz was formulated [20]. This algorithm, which was later generalized to the open chain [21], is restricted to real Bethe roots, requires ancillary qubits, and is probabilistic. Moreover, the success probability was shown to tend super-exponentially to zero with the number of down-spins [22]. A different approach, based instead on the algebraic Bethe ansatz, was subsequently developed for the closed XXZ chain [23]. The latter algorithm is not limited to real Bethe roots, and is deterministic; however, it is not explicit, as it requires performing QR decompositions of matrices whose size scales exponentially with the number of down-spins. Analytic formulae for the unitaries of [23] were proposed in [24].
We present here a new algorithm for preparing arbitrary $L$-qubit $U(1)$-eigenstates, including exact Bethe states of the spin-1/2 XXZ chain with either open or closed boundary conditions. Like the algorithms [20, 21], it is based on the coordinate Bethe ansatz; however, it is not limited to real Bethe roots, does not require ancillary qubits, and it is deterministic. Moreover, unlike [23], our algorithm is explicit, and does not make use of QR decompositions. For a state with $M$ down-spins in a chain of length $L$, the circuit uses only $L$ qubits, and has size (gate count) and depth $\mathcal{O}\left(\binom{L}{M}\right)$. The algorithm makes use of certain complex numbers $f(w)$, which are expensive to compute classically for large $M$. Our algorithm is inspired by an efficient recursive algorithm[25, 26, 27, 28] for preparing Dicke states [29]. Code in Qiskit [30] implementing our algorithm is available as Supplementary Material.
The remainder of this paper is structured as follows. In Section 2, starting from an expression for the Bethe state that we wish to prepare, we define the states (2.4), and observe that they satisfy an elementary recursion (2.7). In Section 3, we use this recursion to derive a corresponding recursion for a unitary operator $U_{L}$ that prepares the Bethe state from a simple product state (3.3), in terms of certain simpler unitary operators $W_{m}$, see (3.7). The latter operators are constructed in terms of yet simpler unitary operators $\textup{I}_{m,l}$, see (3.11). A brief discussion of these results is presented in Section 4. In Appendix A, the coordinate Bethe ansatz for the closed and open spin-1/2 XXZ chains is summarized. Finally, complete circuit diagrams for examples with $(L,M)=(4,2)$ and $(L,M)=(6,3)$ are presented in Appendix B.
@@ 2 Recursion
where the sum is over the set $P(L,M)$ of all permutations $w$ with $L-M$ zeros (up-spins) and $M$ ones (down-spins), which we think of as $\binom{L}{M}$ strings of length $L$. Moreover, the coefficients $f(w)$ are complex numbers that depend on the corresponding $M$ Bethe roots, which we assume are known, see Appendix A. We emphasize that these coefficients are computed classically, with complexity that scales exponentially with $M$. Here $F(\{\})$ denotes the normalization
@@
and $|0011\rangle=|0\rangle\otimes|0\rangle\otimes|1\rangle\otimes|1\rangle$, etc., with computational basis states $|0\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$ and $|1\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$. The Bethe states are not the only examples of states that take the form (2.1); in fact, the eigenstates of any Hamiltonian with $U(1)$ symmetry take this form. [^1] Indeed, since the states $|B^{L}_{M}\rangle$ form an orthonormal basis, any state $|\phi\rangle$ can be expressed as $|\phi\,\rangle=\sum_{M=0}^{L}C_{M}\,|B^{L}_{M}\rangle$. Hence, if $|\phi\rangle$ is an eigenstate of $S^{z}$ with eigenvalue $\alpha$, then $\alpha\,|\phi\rangle=S^{z}|\phi\rangle=\sum_{M=0}^{L}\,(L/2-M)\,C_{M}\,|B^{L}_{M}\rangle$. This implies $\alpha\,C_{M}=(L/2-M)\,C_{M}$, so either $\alpha=L/2-M$ or $C_{M}=0$. In particular, $C_{M}$ is non-zero for exactly one value of $M$, so $|\phi\rangle=|B^{L}_{M}\rangle$. Although we have in mind values $f(w)$ dictated by Bethe ansatz, and we shall refer to $|B^{L}_{M}\rangle$ as Bethe states, our algorithm is actually more general, as it prepares the state $|B^{L}_{M}\rangle$ for arbitrary complex numbers $f(w)$. In particular, for the case $f(w)=1$, $|B^{L}_{M}\rangle$ reduces to the Dicke state $|D^{L}_{M}\rangle$, whose preparation was shown to have an efficient recursive algorithm [25].
@@
That is, we are restricting the sum in (2.4) to permutations $w$ of the form $w=ab$, where $ab$ denotes the concatenation of the strings $a$ and $b$, so $|w\rangle=|a\rangle\otimes|b\rangle$. The state (2.4) is thus the “$b$-tail” of the Bethe state (2.1). In other words, the sum in (2.4) is effectively over $a$, while keeping $b$ fixed. Note that when $b=\{\}$ is the empty string, $F(b)$ reduces to the normalization given by (2.2). For example, for the state $|B^{4}_{2}\rangle$ (2.3) with $b=0$, we have
@@ 3 Algorithm
It follows from the recursion (2.7) that $U_{m}$ has the recursive property
@@
It thus remains to construct a circuit $W_{m}$ that performs (3.5) for all $l$ and $b$. This can be achieved by expressing $W_{m}$ as an ordered product of simpler operators $\textup{I}_{m,l}$ independent of $b$ that perform the job of $W_{m}$ (3.5) for specific values $l$, in such a way that operators with different $l$-values do not interfere with each other. That is, $\textup{I}_{m,l}$ has the following properties:
@@
for all $l$. In view of the bounds (3.2), we conclude that the operator $W_{m}$, independent of $l$, can be written as the following ordered product of I-operators:
@@
Figure 1: Circuit diagrams for $\textup{I}_{m,l}$, with $\mathcal{U}(m,l)$ defined as a product of $u$-gates (3.12). Additional controls on wires 00 to $L-m-1$ (represented by a rectangle with 3 bullets) are placed on each $u(m,l,b)$ at the locations where $b$ contains a $1$. The number of $1$’s in $b$ is $M-l$.
The circuit diagram for an operator $\textup{I}_{m,l}$ with the properties (3.8) and (3.9) is displayed in Fig. 1. Similarly to the Dicke circuit [25, 26, 27, 28], the pair of CNOT gates together with the multi-controlled $\mathcal{U}(m,l)$-gate perform the rotation in the subspace $|10\rangle$ and $|01\rangle$ defined by (3.5). $\mathcal{U}(m,l)$ denotes the product
@@
Each $u$-gate acts on wire $L-m$, and has controls on the wires $L-m+l$ and $L-m+l-1$ (the latter of which is present if and only if $l>1$).444 We follow Qiskit conventions, where e.g. $|110\rangle$ corresponds to $|0\rangle$ on wire 00 and $|1\rangle$ on wires $1$ and $2$. Additionally, because the angles depend on $b$, additional controls are placed on wires 00 to $L-m-1$ at the locations where the string $b$ takes the value $1$, to differentiate between different $b$ [^5] This method has been used before to generate states with many free parameters, see e.g. [31]. It suffices to put controls for the $1$’s in $b$, as opposed to also including negated controls for 00’s. This is because each $\textup{I}_{m,l}$ operator conserves the number of $1$’s, and only acts on states of the form $|0\rangle^{\otimes(m-l)}\,|1\rangle^{\otimes l}\,|b\rangle_{L-m}$, which guarantees that $b$ contains exactly $M-l$ $1$’s. For example, for $L=4,\,M=2$, $u(m=2,l=1,b=01)$ has controls on the wires $3$ and 00, whereas $u(m=2,l=1,b=10)$ has controls on the wires $3$ and $1$. For the case in which $b$ in (3.12) is unique, these additional controls can be omitted. Complete circuit diagrams for examples with $L=4,\,M=2$ and $L=6,\,M=3$ are displayed in Figs. 2 and 3 in Appendix B. An implementation of this algorithm using Qiskit [30] is available as Supplementary Material.
@@
$M+2M(L-M)+{L\choose M}-1=\mathcal{O}\left(\binom{L}{M}\right)\,,$ (3.17)
and the depth is $\mathcal{O}\left(\binom{L}{M}\right)$ as well. For the central case $M=L/2$, we note the asymptotic approximation (see e.g. [32], p. 35)
@@
Further, we see that the number of CNOT-gates (3.15) and $u$-gates (3.16) are symmetric under $M\to L-M$, which implies that there is no advantage to using the dual symmetry $|B^{L}_{M}\rangle=X^{\otimes L}\,|B^{L}_{L-M}\rangle$.
@@ 4 Discussion
We have formulated an algorithm for exactly preparing on a quantum computer an $L$-qubit $U(1)$-eigenstate consisting of an arbitrary superposition of computational basis states with $M$ ones (down-spins) and $L-M$ zeros (up-spins) (2.1). The circuit is ideally suited for models solved by Bethe ansatz, which have available general recipes for the values $f(w)$ corresponding to the models’ eigenstates for arbitrary values of $L$ and $M$ in terms of Bethe roots $\{k_{1},...,k_{M}\}$. However, these coefficients $f(w)$, of which there are (${L\choose M}$ many, are each expressed as a sum over $M!$ terms for the closed chain (A.5), and $2^{M}M!$ terms for the open chain (A.12), and are thus expensive to compute classically for large $M$. (For eigenstates of non-integrable models, the $f(w)$ values are generally not known.) The circuit size is of the same order as the number of free parameters $f(w)$, suggesting that the circuit with generic values of $f(w)$ is optimal for given values of $L$ and $M$. Nevertheless, for $M=L/2$, the circuit size and depth scale exponentially with $L$. As noted in the Introduction, these states could be used to compute correlation functions [17, 18].
@@
We have assumed here, as in [20, 21, 22, 23, 24], that the Bethe roots for the desired state are known. Real Bethe roots can readily be determined classically, even for values $L\,,M\sim 10^{3}$, by iteration (see, e.g. [34] and references therein). However, complex Bethe roots are generally difficult to find classically. Following up on an idea suggested in [19], it may be possible to implement a hybrid classical-quantum approach (see e.g. [7] and references therein) for determining such complex Bethe roots by using our Bethe states as trial states and the unknown Bethe roots as variational parameters. [^6] This approach has now indeed been implemented in [35].
We emphasize that our algorithm makes use of integrability only in so far as computing $f(w)$ (A.5), (A.12), which is performed classically. It would be interesting to find a way to further exploit integrability. We also note that, for given values of $L$ and $M$, the circuit size and depth are independent of the values of $f(w)$. For example, for Dicke states (all $f(w)=1$), our circuit does not reduce to the efficient one in [25]. It would be desirable to find a way to take advantage of simplifications of $f(w)$, which also arise in matrix product state (MPS) representations, the XX chain, etc. While MPS methods give an efficient way to prepare ground states of gapped models [36], implementing these approaches for excited states is more challenging with tensor networks. Further, such methods require an ancillary space that grows with the system size, as well as numerical decompositions of unitaries into elementary gates; this typically makes such algorithms non-explicit.
@@ A.1 Closed periodic XXZ chain
$f(w)=\sum_{\sigma\in S_{M}}\varepsilon(\sigma)\,A(k_{\sigma(1)},\cdots,k_{\sigma(M)})\,e^{i\sum_{j=1}^{M}k_{\sigma(j)}x_{j}}\,,$ (A.5)
where the sum is over the set $S_{M}$ of all permutations $\sigma$ of $\{1,\dots,M\}$, and $\varepsilon(\sigma)=\pm 1$ denotes the sign of the permutation $\sigma$. Moreover,
@@ A.2 Open XXZ chain
$f(w)=\sum_{\sigma\in S_{M}}\sum_{\epsilon_{1},\dots,\epsilon_{M}=\pm 1}\varepsilon(\sigma)\,\epsilon_{1}\dots\epsilon_{M}\,A(\epsilon_{1}k_{\sigma(1)},\ldots,\epsilon_{M}k_{\sigma(M)})\,e^{i\sum_{j=1}^{M}\epsilon_{j}k_{\sigma(j)}x_{j}}\,,$ (A.12)
and now